# Chasing the Tail: Effective Rubric-based Reward Modeling for Large Language Model Post-Training

### v1

@@ 5.2 Reward model accuracy improves in the high-reward tail
Our theoretical analysis (Section˜3) suggests that accuracy in the high-reward tail is the critical factor for downstream RL performance. To understand why refinement with better and more diverse responses helps, we evaluate the agreement between rubric-based rewards and the ground-truth judge, separately on the high- and low-reward regions.
As shown in Table˜2, incorporating any candidate responses through refinement improves rubric accuracy compared to the prompt-only baseline. More importantly, rubrics refined with great pairs largely improve accuracy in the high-reward region, while good pairs improve accuracy more than great pairs in the low-reward region. Iterative refinement with great pairs pushes the accuracy in the high-reward region even further, mirroring the RL improvements in Table˜1. This confirms that both principles work by sharpening reward model accuracy where it matters most: the high-value tail.
@@
Initial 1 Good Pair 1 Great Pair 4 Great Pairs 4 Great & Diverse Pairs
@@ 5.3 Refinements from better responses are more sophisticated
Finally, we analyze how refinements differ when using good versus great candidate responses. To understand how stronger candidate responses lead to better rubrics, we analyzed the types of refinements made when using different quality levels of candidate pairs. We prompted an LLM to compare initial and refined rubrics, and categorized the improvements into semantic clusters (see details in Appendix˜H).
@@
In the example from Appendix˜I, for a medical prompt about a patient with serious symptoms, two initially tied great responses are distinguished by adding the criterion: “The response mentions that urgent imaging (e.g., contrast-enhanced CT or MRI/MRV) is required to confirm the diagnosis.” This refinement, from the “Enhancing verification, validation, and evidence standards” cluster, mandates a critical, verifiable clinical action, and only one of the responses satisfies. Such qualitative results confirm our finding that comparing great responses provides the nuanced distinctions needed to identify excellent outputs, thereby sharpening accuracy in the high-reward tail.
@@ 7 Discussion
While we find a medium amount off-policy responses ($n=5000$, in addition to the same number of on-policy responses) do not help Bradley-Terry reward model guide the current policy (see Appendix˜D), we note that other work successfully train BT reward model with off-policy samples, but with a much larger scale—using up to 20 million high quality samples ((Liu et al., 2025a; Cui et al., 2023)). Indeed, Bradley-Terry reward model’s generalizability scales with the number, and diversity of training samples. However, it’s not always easy to find large-scale data for many specialized domains, such as healthcare. In contrast, rubric-based reward can easily encode generalizable principles from limited amount of data.
@@ Appendix A Theoretical Results
First, we compute the KL divergence. When $f(R_{0}^{x})\sim U(0,1)$, by Proposition˜1, the KL divergence is
[…]
@@
[…]
@@ Appendix B Prompts Used for Experiments
Prompt for Constructing Initial Rubrics You’re a skilled judge evaluating the quality of LLM responses to a user prompt. Your first task is to create a comprehensive rubric for grading these responses across multiple dimensions. Given a user prompt, generate a list of binary (yes/no) criteria. These criteria should assess how well the LLM answered the prompt. Only write rubrics you are confident about. Here are tips for writing good rubrics: i. MECE: - Mutually Exclusive, Collectively Exhaustive ii. Completeness: - Consider all the elements you would want to include to create a perfect response and put them into the rubric. This means including not only the facts and statements directly requested by the prompt, but also the supporting details that provide justification, reasoning, and logic for your response. Each of these elements should have a criterion because each criterion helps to develop the answer to the question from a slightly different angle. iii. No overlapping: - the same error from a model shouldn’t be punished multiple times. iv. Diversity: - The rubric items should include variable types of information. - If all criteria are like “the response mentions A”, “the response mentions B”, then this is not a good rubric. v: How many rubric items for each prompt - There is no golden standard, and the desired number of rubrics varies by accounts and task types. - Write rubrics that cover all aspects of an ideal response. vi: How many rubric items to fail - A good rule of thumb is that the model fails on 50% of rubrics items vii: Atomicity / Non-stacked - Each rubric criterion should evaluate exactly one distinct aspect. Avoid bundling multiple criteria into a single rubric. Most stacked criteria with the word “and” can be broken up into multiple pieces. Response identifies George Washington as the first U.S. president and mentions he served two terms. Response identifies George Washington as the first U.S. president. Response mentions that George Washington served two terms. viii: Specificity - Criteria should be binary (true or false) and objective. - Avoid vague descriptions (e.g., "the response must be accurate" is vague). - Example: "The response should list exactly three examples." ix: Self-contained - Each criterion should contain all the information needed to evaluate a response, e.g. Mentions the capital city of Canada. Mentions the capital city of Canada is Ottawa. x: Criterion should be verifiable without requiring external search. Response names any of the Nobel Prize winners in Physics in 2023 Response names any of the following Nobel Prize winners in Physics in 2023: Pierre Agostini, Ferenc Krausz, or Anne L’Huillier. xi. The binary criteria should be phrased so that yes means the model response is good and no means the model response is bad. Finally, we want to assign different weight for each question. Give a weight on a scale of 1 (least important) to 3 (most important) for each question based on 1. the question’s alignment with user demand (3 if user would be frustrated if the answer is no; 1 if user would not be bothered at all if the answer is no) 2. the question’s importance in terms of determining quality/correctness (3 if the response would be completely incorrect if the answer is no; 1 if an extreme edge case would be missed and the overall quality won’t be affected if the answer is no) Here is the user prompt for which we want to generate a rubric: PROMPT: {prompt} Return ONLY the JSON array of the rubrics, no other text. For example: Note: Local IDs will be automatically assigned to each criterion (c1, c2, c3, etc.), so don’t include IDs into outputed criterion.
@@
Prompt for Improving Rubrics You’re a skilled judge assessing the quality of LLM responses to a user prompt. The current rubric isn’t good enough to effectively differentiate between high-quality responses. Your goal is to improve the current rurbics to address this (adding new creteria, rewriting, decomposing, and deleting the current creteria). The updated rubric must be comprehensive and consistently applicable for grading LLM responses. These criteria should specifically assess how well the LLM answered the given prompt. Only write rubrics you are confident about. Here are tips for writing good rubrics: i. MECE: - Mutually Exclusive, Collectively Exhaustive ii. Completeness: - Consider all the elements you would want to include to create a perfect response and put them into the rubric. This means including not only the facts and statements directly requested by the prompt, but also the supporting details that provide justification, reasoning, and logic for your response. Each of these elements should have a criterion because each criterion helps to develop the answer to the question from a slightly different angle. iii. No overlapping: - the same error from a model shouldn’t be punished multiple times. iv. Diversity: - The rubric items should include variable types of information. - If all criteria are like “the response mentions A”, “the response mentions B”, then this is not a good rubric. v: How many rubric items for each prompt - There is no golden standard, and the desired number of rubrics varies by accounts and task types. - Write rubrics that cover all aspects of an ideal response. vi: How many rubric items to fail - A good rule of thumb is that the model fails on 50% of rubrics items vii: Atomicity / Non-stacked - Each rubric criterion should evaluate exactly one distinct aspect. Avoid bundling multiple criteria into a single rubric. Most stacked criteria with the word “and” can be broken up into multiple pieces. Response identifies George Washington as the first U.S. president and mentions he served two terms. Response identifies George Washington as the first U.S. president. Response mentions that George Washington served two terms. viii: Specificity - Criteria should be binary (true or false) and objective. - Avoid vague descriptions (e.g., "the response must be accurate" is vague). - Example: "The response should list exactly three examples." ix: Self-contained - Each criterion should contain all the information needed to evaluate a response, e.g. Mentions the capital city of Canada. Mentions the capital city of Canada is Ottawa. x: Criterion should be verifiable without requiring external search. Response names any of the Nobel Prize winners in Physics in 2023 Response names any of the following Nobel Prize winners in Physics in 2023: Pierre Agostini, Ferenc Krausz, or Anne L’Huillier. xi. The binary criteria should be phrased so that yes means the model response is good and no means the model response is bad. Finally, we want to assign different weight for each criterion. Give a weight on a scale of 1 (least important) to 3 (most important) for each question based on 1. the question’s alignment with user demand (3 if user would be frustrated if the answer is no; 1 if user would not be bothered at all if the answer is no) 2. the question’s importance in terms of determining quality/correctness (3 if the response would be completely incorrect if the answer is no; 1 if an extreme edge case would be missed and the overall quality won’t be affected if the answer is no) Here is the user prompt for which we want to improve the rubric: PROMPT: {prompt} The existing rubrics we are using is: {rubrics} The two reference responses are: Reponse 1: {response1} Reponse 2: {response2} Return ONLY the JSON array of the full rubrics, no other text. For example: Note: Local IDs will be automatically assigned to each criterion, so don’t include IDs in your output.
@@
Prompt for Scoring Responses You are a skilled judge who will be assessing the quality of LLM responses to a user prompt. Given a user prompt, LLM response, and a rubric, your task is evalauting the performance of the model response by seeing whether or not it meets the rubric dimension. Answer the each of the given rubric dimension in either "yes" or "no". Do not output any response other than "yes" or "no". Keep in mind that you will be grading industry-leading LLMs. Make sure to have high expectation for grading the responses. Make sure your evaluation is as objective and consistent as it could be. By consistent we mean that a different evaluator’s assessment of the task should agree with yours. Think carefully before you make the decision. After you make the decision, explicitly output which dimension receives "yes" and which dimension receives "no". Input: * PROMPT: {prompt} * RESPONSE: {response} * RUBRIC: {rubric} Return ONLY the JSON array, no other text. For example:
@@ Appendix C Hyperparameter
The hyperparameter for GRPO training of RLRR is presented in Table˜4.
@@ Appendix D Empirical Results on RLHF
We finetune a Bradley-Terry Reward model on various responses, with preference generated by GPT-4.1, the same model as the judge model for evaluation. For each of the prompt in the training set, we generated a pair of responses at temperature 1.0 using the base policy model Qwen3-8B-Base (on-policy) or Gemini-2.5-Pro (off-policy). Preferences were labeled using GPT-4.1, the same model used for final evaluation. This preference data was then used to train a reward model based on Llama-3.1-8B-Instruct, with hyperparameters specified in Table˜6. Finally, this reward model was used for GRPO training, following the configuration in Table˜4.
@@ Appendix H Pattern detection on rubric refinements
In Table˜3 we show clusters of various rubric-refinements. To get this result, we first prompt GPT-4.1 to describe the differences between refined rubrics, and the initial rubrics (for refinement with one-good-pair, and one-great-pair). The differences are listed as atomic items describing different aspects of improvement. We then run a clustering algorithm on the atomic differences from both one-good-pair, and one-great-pair (they have almost the same number of atomic differences) to get interpretable clusters, and report the proportions from one-good-pair versus and one-great-pair
@@ Appendix I Examples of Rubrics and Rubric Refinements
In this example, the initial rubric cannot differentiate between the two great responses (both from Gemini-2.5-Pro). This is because it primarily handles basic criteria (e.g. the diagnosis is correct), which both responses satisfy. The refinement process resolves the tie by adding a new criterion c7—requiring the mention of an imaging test to confirm the diagnosis. Such refinement is one example for the cluster Enhancing verification, validation, and evidence standards identified in Table˜3.
@@
Criterion (Refined) Weight R1 (0/1) R2 (0/1) c1: Identifies CST as most likely diagnosis 3 1 1 c2: Explicitly states CST is a medical emergency 3 1 1 c3: Links medial canthus pain + decongestants to sinusitis/dacryocystitis 3 1 1 c4: Diplopia due to abducens (CN VI) involvement 3 1 1 c5: Papilledema from impaired venous drainage/ICP 2 1 1 c6: Sepsis secondary to CST (chills/rigors) 2 1 1 c7: Urgent imaging (contrast CT or MRI/MRV) required to confirm 2 1 0     diagnosis c8: High-dose IV antibiotics are initial mainstay 2 1 1 c9: Medical disclaimer / seek immediate care 2 0 0 c10: Mentions orbital cellulitis differential 1 1 1 c11: Notes other CNs (III, IV, V1, V2) may be affected 1 1 0 c12: Avoids incorrect primary diagnosis 3 1 1

### v2

@@ 5.2 Reward model accuracy improves in the high-reward tail
Our theoretical analysis (Section 3) suggests that accuracy in the high-reward tail is the critical factor for downstream RL performance. To understand why refinement with better and more diverse responses helps, we evaluate the agreement between rubric-based rewards and the ground-truth judge, separately on the high- and low-reward regions.
As shown in Table 2, incorporating any candidate responses through refinement improves rubric accuracy compared to the prompt-only baseline. More importantly, rubrics refined with great pairs largely improve accuracy in the high-reward region, while good pairs improve accuracy more than great pairs in the low-reward region. Iterative refinement with great pairs pushes the accuracy in the high-reward region even further, mirroring the RL improvements in Table 1. This confirms that both principles work by sharpening reward model accuracy where it matters most: the high-value tail.
@@
Initial Only 1 Good Pair 1 Great Pair 4 Great Pairs 4 Great & Diverse Pairs
@@ 5.3 Refinements from better responses are more sophisticated
Finally, we analyze how refinements differ when using good versus great candidate responses. To understand how stronger candidate responses lead to better rubrics, we analyzed the types of refinements made when using different quality levels of candidate pairs. We prompted an LLM to compare initial and refined rubrics, and categorized the improvements into semantic clusters (see details in Appendix H).
@@
In the example from Appendix I, for a medical prompt about a patient with serious symptoms, two initially tied great responses are distinguished by adding the criterion: “The response mentions that urgent imaging (e.g., contrast-enhanced CT or MRI/MRV) is required to confirm the diagnosis.” This refinement, from the “Enhancing verification, validation, and evidence standards” cluster, mandates a critical, verifiable clinical action, and only one of the responses satisfies. Such qualitative results confirm our finding that comparing great responses provides the nuanced distinctions needed to identify excellent outputs, thereby sharpening accuracy in the high-reward tail.
@@ 7 Discussion
While we find a medium amount off-policy responses ($n=5000$, in addition to the same number of on-policy responses) do not help Bradley-Terry reward model guide the current policy (see Appendix D), we note that other work successfully train BT reward model with off-policy samples, but with a much larger scale—using up to 20 million high quality samples ((Liu et al., 2025a; Cui et al., 2023)). Indeed, Bradley-Terry reward model’s generalizability scales with the number, and diversity of training samples. However, it’s not always easy to find large-scale data for many specialized domains, such as healthcare. In contrast, rubric-based reward can easily encode generalizable principles from limited amount of data.
@@ Appendix A Theoretical Results
First, we compute the KL divergence. When $f(R_{0}^{x})\sim U(0,1)$, by Proposition 1, the KL divergence is
$\mathbb{D}_{\mathrm{KL}}\left[\pi_{r}(y\mid x)\|\pi_{0}(y\mid x)\right]=\mathbb{E}_{x\sim D}\left[\frac{\mathbb{E}\left[f(R^{x}_{0})~e^{f(R^{x}_{0})/\beta}/{\beta}\right]}{\mathbb{E}\left[e^{f(R^{x}_{0})/\beta}\right]}-\log\mathbb{E}\left[e^{f(R^{x}_{0})/\beta}\right]\right]$
$=\mathbb{E}_{x\sim D}\left[\frac{\int_{0}^{1}ue^{u/\beta}\mathrm{d}u}{\beta\int_{0}^{1}e^{u/\beta}\mathrm{d}u}-\log\left(\int_{0}^{1}e^{u/\beta}\mathrm{d}u\right)\right]=\frac{(1/\beta-1)e^{1/\beta}+1}{e^{1/\beta}-1}-\log\left[\beta(e^{1/\beta}-1)\right].$
@@
$\mathbb{E}_{x\sim D,~y\sim\pi_{r}(\cdot\mid x)}\left[r^{\star}(x,y)\right]=\mathbb{E}_{x\sim D}\left[\frac{\mathbb{E}\left[R^{x}_{0}~e^{f(R^{x}_{0})/\beta}\right]}{\mathbb{E}\left[e^{f(R^{x}_{0})/\beta}\right]}\right]$
$=\mathbb{E}_{x\sim D}\left[\frac{\mathbb{E}\left[f^{-1}(T_{0}^{x})~e^{T^{x}_{0}/\beta}\right]}{\mathbb{E}\left[e^{T_{0}^{x}/\beta}\right]}\right]=\mathbb{E}_{x\sim D}\left[\frac{\int_{0}^{1}f^{-1}(u)e^{u/\beta}\mathrm{d}u}{\int_{0}^{1}e^{u/\beta}\mathrm{d}u}\right]$
$=\frac{\int_{0}^{1}f^{-1}(u)e^{u/\beta}\mathrm{d}u}{\int_{0}^{1}e^{u/\beta}\mathrm{d}u}=\frac{\int_{0}^{1}f^{-1}(u)e^{u/\beta}\mathrm{d}u}{\beta(e^{u/\beta}-1)}$
@@ Appendix B Prompts Used for Experiments
Prompt for Constructing Initial Rubrics You’re a skilled judge evaluating the quality of LLM responses to a user prompt. Your first task is to create a comprehensive rubric for grading these responses across multiple dimensions. Given a user prompt, generate a list of binary (yes/no) criteria. These criteria should assess how well the LLM answered the prompt. Only write rubrics you are confident about. Here are tips for writing good rubrics: i. MECE: - Mutually Exclusive, Collectively Exhaustive ii. Completeness: - Consider all the elements you would want to include to create a perfect response and put them into the rubric. This means including not only the facts and statements directly requested by the prompt, but also the supporting details that provide justification, reasoning, and logic for your response. Each of these elements should have a criterion because each criterion helps to develop the answer to the question from a slightly different angle. iii. No overlapping: - the same error from a model shouldn’t be punished multiple times. iv. Diversity: - The rubric items should include variable types of information. - If all criteria are like “the response mentions A”, “the response mentions B”, then this is not a good rubric. v: How many rubric items for each prompt - There is no golden standard, and the desired number of rubrics varies by accounts and task types. - Write rubrics that cover all aspects of an ideal response. vi: How many rubric items to fail - A good rule of thumb is that the model fails on 50% of rubrics items vii: Atomicity / Non-stacked - Each rubric criterion should evaluate exactly one distinct aspect. Avoid bundling multiple criteria into a single rubric. Most stacked criteria with the word “and” can be broken up into multiple pieces. Response identifies George Washington as the first U.S. president and mentions he served two terms. Response identifies George Washington as the first U.S. president. Response mentions that George Washington served two terms. viii: Specificity - Criteria should be binary (true or false) and objective. - Avoid vague descriptions (e.g., ”the response must be accurate” is vague). - Example: ”The response should list exactly three examples.” ix: Self-contained - Each criterion should contain all the information needed to evaluate a response, e.g. Mentions the capital city of Canada. Mentions the capital city of Canada is Ottawa. x: Criterion should be verifiable without requiring external search. Response names any of the Nobel Prize winners in Physics in 2023 Response names any of the following Nobel Prize winners in Physics in 2023: Pierre Agostini, Ferenc Krausz, or Anne L’Huillier. xi. The binary criteria should be phrased so that yes means the model response is good and no means the model response is bad. Finally, we want to assign different weight for each question. Give a weight on a scale of 1 (least important) to 3 (most important) for each question based on 1. the question’s alignment with user demand (3 if user would be frustrated if the answer is no; 1 if user would not be bothered at all if the answer is no) 2. the question’s importance in terms of determining quality/correctness (3 if the response would be completely incorrect if the answer is no; 1 if an extreme edge case would be missed and the overall quality won’t be affected if the answer is no) Here is the user prompt for which we want to generate a rubric: PROMPT: {prompt} Return ONLY the JSON array of the rubrics, no other text. For example: Note: Local IDs will be automatically assigned to each criterion (c1, c2, c3, etc.), so don’t include IDs into outputed criterion.
@@
Prompt for Improving Rubrics You’re a skilled judge assessing the quality of LLM responses to a user prompt. The current rubric isn’t good enough to effectively differentiate between high-quality responses. Your goal is to improve the current rurbics to address this (adding new creteria, rewriting, decomposing, and deleting the current creteria). The updated rubric must be comprehensive and consistently applicable for grading LLM responses. These criteria should specifically assess how well the LLM answered the given prompt. Only write rubrics you are confident about. Here are tips for writing good rubrics: i. MECE: - Mutually Exclusive, Collectively Exhaustive ii. Completeness: - Consider all the elements you would want to include to create a perfect response and put them into the rubric. This means including not only the facts and statements directly requested by the prompt, but also the supporting details that provide justification, reasoning, and logic for your response. Each of these elements should have a criterion because each criterion helps to develop the answer to the question from a slightly different angle. iii. No overlapping: - the same error from a model shouldn’t be punished multiple times. iv. Diversity: - The rubric items should include variable types of information. - If all criteria are like “the response mentions A”, “the response mentions B”, then this is not a good rubric. v: How many rubric items for each prompt - There is no golden standard, and the desired number of rubrics varies by accounts and task types. - Write rubrics that cover all aspects of an ideal response. vi: How many rubric items to fail - A good rule of thumb is that the model fails on 50% of rubrics items vii: Atomicity / Non-stacked - Each rubric criterion should evaluate exactly one distinct aspect. Avoid bundling multiple criteria into a single rubric. Most stacked criteria with the word “and” can be broken up into multiple pieces. Response identifies George Washington as the first U.S. president and mentions he served two terms. Response identifies George Washington as the first U.S. president. Response mentions that George Washington served two terms. viii: Specificity - Criteria should be binary (true or false) and objective. - Avoid vague descriptions (e.g., ”the response must be accurate” is vague). - Example: ”The response should list exactly three examples.” ix: Self-contained - Each criterion should contain all the information needed to evaluate a response, e.g. Mentions the capital city of Canada. Mentions the capital city of Canada is Ottawa. x: Criterion should be verifiable without requiring external search. Response names any of the Nobel Prize winners in Physics in 2023 Response names any of the following Nobel Prize winners in Physics in 2023: Pierre Agostini, Ferenc Krausz, or Anne L’Huillier. xi. The binary criteria should be phrased so that yes means the model response is good and no means the model response is bad. Finally, we want to assign different weight for each criterion. Give a weight on a scale of 1 (least important) to 3 (most important) for each question based on 1. the question’s alignment with user demand (3 if user would be frustrated if the answer is no; 1 if user would not be bothered at all if the answer is no) 2. the question’s importance in terms of determining quality/correctness (3 if the response would be completely incorrect if the answer is no; 1 if an extreme edge case would be missed and the overall quality won’t be affected if the answer is no) Here is the user prompt for which we want to improve the rubric: PROMPT: {prompt} The existing rubrics we are using is: {rubrics} The two reference responses are: Reponse 1: {response1} Reponse 2: {response2} Return ONLY the JSON array of the full rubrics, no other text. For example: Note: Local IDs will be automatically assigned to each criterion, so don’t include IDs in your output.
@@
Prompt for Scoring Responses You are a skilled judge who will be assessing the quality of LLM responses to a user prompt. Given a user prompt, LLM response, and a rubric, your task is evalauting the performance of the model response by seeing whether or not it meets the rubric dimension. Answer the each of the given rubric dimension in either ”yes” or ”no”. Do not output any response other than ”yes” or ”no”. Keep in mind that you will be grading industry-leading LLMs. Make sure to have high expectation for grading the responses. Make sure your evaluation is as objective and consistent as it could be. By consistent we mean that a different evaluator’s assessment of the task should agree with yours. Think carefully before you make the decision. After you make the decision, explicitly output which dimension receives ”yes” and which dimension receives ”no”. Input: * PROMPT: {prompt} * RESPONSE: {response} * RUBRIC: {rubric} Return ONLY the JSON array, no other text. For example:
@@ Appendix C Hyperparameter
The hyperparameter for GRPO training of RLRR is presented in Table 4. The finance domain is an exception due to the limited amount of available data, and its hyperparameters are presented separately in Table 5.
@@
Table 5: GRPO Hyperparameter Configuration in Finance Domain
Hyperparameter Value
Rollouts per Prompt 16
Gradient Accumulation Steps 2
Per-Device Train Batch Size 6
Warmup Ratio 0.1
KL Coefficient 0.01
Learning Rate $1.0\times 10^{-5}$
Learning Rate Scheduler Constant with Warmup
Maximum Sequence Length 3584
Training Steps 320
@@ Appendix D Empirical Results on RLHF
We finetune a Bradley-Terry Reward model on various responses, with preference generated by GPT-4.1, the same model as the judge model for evaluation. For each of the prompt in the training set, we generated a pair of responses at temperature 1.0 using the base policy model Qwen3-8B-Base (on-policy) or Gemini-2.5-Pro (off-policy). Preferences were labeled using GPT-4.1, the same model used for final evaluation. This preference data was then used to train a reward model based on Llama-3.1-8B-Instruct, with hyperparameters specified in Table 7. Finally, this reward model was used for GRPO training, following the configuration in Table 4.
@@ Appendix H Pattern detection on rubric refinements
In Table 3 we show clusters of various rubric-refinements. To get this result, we first prompt GPT-4.1 to describe the differences between refined rubrics, and the initial rubrics (for refinement with one-good-pair, and one-great-pair). The differences are listed as atomic items describing different aspects of improvement. We then run a clustering algorithm on the atomic differences from both one-good-pair, and one-great-pair (they have almost the same number of atomic differences) to get interpretable clusters, and report the proportions from one-good-pair versus and one-great-pair
@@ Appendix I Examples of Rubrics and Rubric Refinements
In this example, the initial rubric cannot differentiate between the two great responses (both from Gemini-2.5-Pro). This is because it primarily handles basic criteria (e.g. the diagnosis is correct), which both responses satisfy. The refinement process resolves the tie by adding a new criterion c7—requiring the mention of an imaging test to confirm the diagnosis. Such refinement is one example for the cluster Enhancing verification, validation, and evidence standards identified in Table 3.
@@
Criterion (Refined) Weight R1 (0/1) R2 (0/1) c1: Identifies CST as most likely diagnosis 3 1 1 c2: Explicitly states CST is a medical emergency 3 1 1 c3: Links medial canthus pain + decongestants to sinusitis/dacryocystitis 3 1 1 c4: Diplopia due to abducens (CN VI) involvement 3 1 1 c5: Papilledema from impaired venous drainage/ICP 2 1 1 c6: Sepsis secondary to CST (chills/rigors) 2 1 1 c7: Urgent imaging (contrast CT or MRI/MRV) required to confirm diagnosis 2 1 0 c8: High-dose IV antibiotics are initial mainstay 2 1 1 c9: Medical disclaimer / seek immediate care 2 0 0 c10: Mentions orbital cellulitis differential 1 1 1 c11: Notes other CNs (III, IV, V1, V2) may be affected 1 1 0 c12: Avoids incorrect primary diagnosis 3 1 1